Suppose you went short with a 940 Mar 20 soybean put. The premium on this put was 25’3. Suppose you were exercised on when the Mar 20 soybean futures were trading at 910.

  1. What futures position were given and at what price were you given that futures position?
  2. How much was the intrinsic value on the 940 Mar 20 soybean put when you were exercised on?  
  3. How much did you make or lose on your short 940 Mar 20 soybean put?
Answer #1

answer.

a. Going for a put short is a position for bullish market in future.The strike price was 940 and premium received by trader was 25.3. Therefore, trader does not see market lower than 940-25.3= 914.7

b. Strike Price= 940 Spot Price closing= 910 Intrinsic value of put= Strike - Spot = 940-910 = 30

C. Premium Received by selling put = 25.3

Value of Put to be paid on exercise = 30

Net payoff = 25.3 - 30 = -4.7

The trader made a loss of 4.7dollars.